Use the Pythagorean theorem to determine the length of each line segment. Where appropriate, round to the nearest hundredth.
The plotted points are at (1, -1) and (3, 2).

I am not understanding this problem and need some help with this. Thanks

Answer by checkley75(3666)

To define the right triangle we need to calculate the x & y distances. Thus:
The x distance is 3-1=2
The y distance is 2-(-1), 2+1=3
So we have the formula x^2+y^2=h^2 h being the hypotenuse.
2^2+3^2=h^2
4+9=h^2
13=h^2 now calculate the sqrt of 13 & h.
3.6=h is the hypotenuse or the distance between these two points.
To become familiar with this concept use graph paper, plot the two points, draw the triangle and measure the hypotenuse (distance between these two points).